OpenStudy (dmezzullo):

Solve for x: 1 over 4 (2x − 14) = 4 x = 15 x = 15 over 2 x = 9 x = 11

OpenStudy (anonymous):

x=15

OpenStudy (dmezzullo):

SHOW WORK

OpenStudy (jiteshmeghwal9):

@dmezzullo typing same questions again & again ?? are u nt getting the concept ??

OpenStudy (anonymous):

I CANT! ill just stop tellin u, u can figure it out

OpenStudy (jiteshmeghwal9):

similar*

OpenStudy (dmezzullo):

@mckenziemichelle im sorry no keep helping

OpenStudy (anonymous):

Na.

OpenStudy (dmezzullo):

:'(

OpenStudy (anonymous):

Distribute. Then solve for x.

OpenStudy (jiteshmeghwal9):

if u want to cry make sign T_T

OpenStudy (anonymous):

Sorry..i'm just not doin anything right so forget it.

OpenStudy (dmezzullo):

yesh u r im just asking if someone can show me the work @mckenziemichelle

OpenStudy (anonymous):

okay

OpenStudy (dmezzullo):

no

OpenStudy (dmezzullo):

omg im sorry im so stupid

OpenStudy (dmezzullo):

i divided 224/32 and got 7.03125

OpenStudy (dmezzullo):

Solve for x: 1 over 4 (2x − 14) = 4 x = 15 x = 15 over 2 x = 9 x = 11

OpenStudy (dmezzullo):

nope.... parentheses r next to the division bar

OpenStudy (jiteshmeghwal9):

\[\LARGE{\frac{1}{4}(2x-14)=4}\]\[\LARGE{\frac{2x-14}{4}=4}\]by cross multiplication.we have,\[\LARGE{2x-14=16}\]\[\LARGE{2x=16+14}\]\[\LARGE{2x=30}\]\[\LARGE{x=15}\]

OpenStudy (dmezzullo):

oh ok

OpenStudy (dmezzullo):

tyvm

OpenStudy (jiteshmeghwal9):

yw :)